I need help setting this problem up.
Tina,Dawn and Harry have $202.50 together. Tina has 3 times as much $ as Dawn. Dawn has twice as much $ as Harry. How much $ does each one have?

Answer by mananth(16946)

Harry has $x
Dawn has $2x
Tina has 3*(2x)=6x
Total =202.50
x+2x+6x=202.50
9x=202.50
/9
x=22.50
Harry has $22.5
Dawn has 2*22.5=$45
Tina has 6*22.5 = $ 135.50